When N2 is converted into NH3 , the equivalent weight of nitrogen will be

Grade:11

4 Answers

Gopal
6 years ago
Oxidation number of N in N2= Zero
Oidation number of N in NH3= -3
Change in oidation number is by 3 units for each N atom , so for N2 molecule it is by 6 units
Therefore equivalent weight =M/6 =28/6
Divyanshu Kharwar
5 years ago
Simple Method-->
Oxidation no. Of N2 = 14
 Oxidation State vary from (N2)zero to (NH3)-3
(neglect - sign).
14/3 = 4.67
Divyanshu Kharwar
5 years ago
Error in my above given answer is in place of oxidation number that is molecular mass of N2. Soory (Unable to edit above answer).
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem below.
 
Oxidation number of N in N2 = Zero
Oxidation number of N in NH3= -3
Change in oidation number is by 3 units for each N atom , so for N2 molecule it is by 6 units
Therefore equivalent weight of N2 = Molecular mass of N2/n factor = M/6 = 28/6 = 4.67 g
 
Thanks and regards,
Kushagra